Circular No. 5747

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V1974 CYGNI S. Shore, Goddard High Resolution Spectrograph (GHRS) Science

Team; S. Starrfield, Arizona State University; G. Sonneborn, Goddard Space Flight Center; T. Ake, GHRS Science Team; and C. Gonzalez-Riestra, IUE Vilspa, report: “On Apr. 1 we obtained a second GHRS observation of V1974 Cyg = Nova Cyg 1992. The data are all mediumresolution (R = 15 000), large-aperture, 20-min integrations. We have observed the following lines: N V 124.0, O IV]/S IV] 140.0, N IV] 148.6, C IV 155.0, [Ne V] 157.5, [Ne IV] 160.2, He II 164.0, O III] 166.5, N III] 175.0 and C III] 191.0 nm. The preliminary results show that the emissionline profiles are now all optically thin. In our earlier observations on 1992 Sept. 7 we found that the resonance lines were all still optically thick. The [Ne IV], [Ne V], He II, and N IV] profiles are essentially identical, showing strong (50 percent or greater) contrast between the emission knots and the diffuse emission. The line widths are also the same, FWHM of ∼ 2000 km/s (the wings being slightly narrower than seen in the earlier spectra). The C IV and N V profiles are quite similar, also displaying many knots of emission (much like those seen on Mg II 280.0 nm since ∼ 1992 May 20). An important result of this new observation is the confirmation that Lyα was in emission last September, while it is now absent from the spectrum: this supports its identification in the low-resolution IUE data. The ejecta appear now to be spectroscopically dominated by a relatively small number of strong emission filaments, moving at a comparatively low velocity (∼ 1000 km/s). We are continuing monitoring with IUE. Any observers who are able to obtain groundbased spectrophotometry are urged to e-mail their results to one of the first authors (hrsshore@stars.gsfc.nasa.gov or starrfie@hydro.la.asu.edu). We are especially interested in He II 468.6 and [O II] and [O III] profiles at comparable resolution to the GHRS spectra.”

Visual magnitude estimates by B. H. Granslo, Fjellhamar, Norway: Mar. 17.01 UT, 11.0; 20.05, 11.2; 23.95, 11.2; 27.85, 11.2; 29.93, 11.3.

SUPERNOVA 1993J IN NGC 3031 Photoelectric photometry: Apr. 2.49 UT, V = 11.50 (T. Kato, Kyoto

University, referred to Corwin’s data for GSC 4383.0928); 2.62, B = 11.78 :, I = 11.08 : (Kato); 2.91, V = 11.36±0.13, B −V = +0.50±0.19 (D. Hanzl, Brno); 3.56, V = 11.76, B = 12.09, I = 11.15 (Kato).

1993 April 5 Brian G. Marsden